In youe math class, 60% of the students are girls. If there are 15 girls in the class, how many students are in your math class?
Answer by stanbon(75887)

In youe math class, 60% of the students are girls. If there are 15 girls in the class, how many students are in your math class?
----
Let the number in the class be "x"::
Equation:
# of girls = # of girls
----------------------
0.60x = 15
x = 15/0.6 = 150/6 = 25 (# in the class)
----------------
Cheers,
Stan H.
--------------